Math 291 Lab 6 Due Monday February 26, 2018

Instructions: Turn in your lab by emailing it to fagerstrom@mnstate.edu. You should email both your raw TEX (.tex) file and your compiled document (in .pdf form).

You will be graded on both your raw TEX code and the accuracy of your compiled document. Name your file in the format of JaneDoeLab2.tex

Create a document that has each of the following elements.

• It is created in L^{A}TEX using the beamer document class.

• Add the necessary information to create a title page for your your beamer presentation, including at least your name, the phrase “Math 291 L^{A}TEX”, a title, and the institution Minnesota State University Moorhead.

• Either pick one of the prepackaged beamer styles and use it for your beamer presentation or design your own. If you design your own, specify an inner, outer, font, and color theme (see the Beamer User Guide for details on and options for these commands).

• Add the commands necessary to add a table of contents to your presentation.

• Use at least three sections in your presentation.

• Each slide should have a title.

• The first content-based frame should contain a reasonably complicated mathematical formula.

• The second content-based frame should use the pause command at least three times.

• The third content-based frame should have an itemized list.

• The fourth content-based frame should make use of the “only” command (with at least four different phrases/slides being created with that command).

• The fifth content-based frame should be identical in content to the fourth content-based frame, but should use the “onslide” command instead of the “only” command.

• The sixth content-based frame should make use of an enumerated list, overlay arguments, and a transparency effect.

Note: Other than these restrictions, you may design the content of your presentation however you wish. You may also add additional frames if you wish, just make sure that you include the above requirements in the specified order.

Note2: If you are using a modified presentation from another of your classes for this assignment, or intend to use this Math 291 assignment as a first step towards an assignment in another class, tell me which class it is. [Doing so is fine, I just want to be informed of the overlap between classes.]

Note3: Remember that the debugging lab is also due on Monday Feb. 26.